OpenStudy (anonymous):

a man who weighs 126 on earth weighs onloy 21 pounds on earth. how much would a person who weighs 25 pounds on the moon weigh on earth

OpenStudy (anonymous):

Do you mean only 21 pounds on moon?

OpenStudy (anonymous):

126/21 = x/25 x = 150 150 pounds on earth